a wire has a diameter of 0.050 inches its cross-sectional area in circular miles is?
1
Expert's answer
2013-09-04T09:07:29-0400
Area= PI * R^2 = PI * (d / 2)^2 = PI * d^2 / 4 = 3.14 * 0.05^2 / 4 = 0.0019625 (sqr inches) = 2498.73 circular mils.